# Learning to crawl: benefits and limits of centralized vs distributed control

### v2

@@ I Introduction
Cephalopods, such as octopus, squid and cuttlefish, have evolved a complex nervous system whose size is comparable to that of vertebrates and endows them with impressive cognitive capabilities. They do have a central brain; however most of their nervous system is distributed throughout their arms which are covered in hundreds of suckers that can sense the environment as well as adhere to substrates. Like all soft-bodied animals, these organisms have to coordinate many degrees of freedom to achieve locomotion. For example, octopuses translocate on surfaces by leveraging their taste-by-touch sense, whereby suckers adhere to surfaces and sense them chemically upon contact [8]. The sensory information acquired by the periphery guides suckers’ adhesion, yielding displacement of the organism along the surface. However, to what extent locomotion is controlled by the central nervous system vs by the ganglia distributed along the arms is not known [9, 10]. Intriguingly, coordination may or may not depend on precise proprioceptive feedback, since the molecular mechanisms for proprioception in cephalopods are unknown [11, 12].
@@
We draw inspiration from these natural crawlers and focus on a toy model to ask what are the pros and cons of centralization vs distribution. The crawler is modeled as a collection of suckers distributed along a straight line and connected by springs, and endowed with a rudimentary form of proprioception (extension vs elongation of the springs) and a rudimentary form of control of the suckers (adhere vs not adhere). Contraction of the springs is not controlled by the crawler but rather follows a stereotyped wave established by a central pattern generator (CPG). If suckers were to adhere at random times, the crawler center of mass would be unable to move on average: is it possible to learn a pattern of adhesion that enables net translocation?
@@
To answer this question we develop a simulator of the crawler and let it learn to translocate by trial and error using the Q-learning Reinforcement Learning algorithm. We compare the learnability, performance and robustness of centralized vs distributed learning architectures. Despite the rudimentary proprioception and control, we find that the sole existence of a common reward –here the speed of the center of mass– is sufficient for independent suckers to learn crawling. When all suckers are simultaneously controlled by a single agent, crawling is both faster and more robust to failure, partially overcoming the crawler’s poor sensorimotor ability. However, the computational cost of the central control scales exponentially with the number of suckers and springs. We then introduce a shallow hierarchical architecture where control is centralized in few control centers. We find that this partial centralization achieves nearly optimal performance and robustness with a limited increase in computational burden. These results complement previous work that focus on controlling contraction rather than adhesion with either purely centralized [13] or purely distributed [14] architectures as well as works where crawling emerges from prescribed control patterns, rather than learning (see e.g. [15, 6])
@@
We hope that our work will serve as a benchmark to explore the optimal level of centralization in more realistic scenarios for natural and artificial crawling, where constraints on efficiency, robustness and computational costs will all concur to define optimality.
@@ II Model and learning
Figure 1: a) Sketch of the 1D crawler model. b) Sketch of the distributed architecture, including the 4 states available to each agent (individual sucker), the 2 actions the sucker controls and the 4$\times$2 Q-matrix for each sucker. c) Centralized architecture: one or more Control Center (CC) control a subset of contiguous suckers. The states for each CC correspond to all combinations of compression/elongation states of the springs within the CC; the actions are all combinations of adhesion/no-adhesion of suckers within the CC. For multiagency (whether the individual agents are suckers or CCs), the “hive update” corresponds to forcing all agents to agree, i.e. write and read the same $Q$ matrix.
@@ II.1 The 1D crawler model
The model, illustrated in fig. 1, is strongly inspired by the one proposed in [15] and considers the crawler as a series of blocks connected by springs along the body axis. The crawler is placed on a flat substrate and its dynamics is dictated by the springs and the interaction with the substrate. We assume friction is proportional to the sliding velocity of the blocks, akin to a viscous term. Each block can actively modify its viscous coefficient thus playing the role of a sucker. In most of the results we assume the friction coefficients are either 0 or infinite. The blocks move via an elongation-contraction wave that propagates from the tail to the head of the crawler. We assume that the CPG is controlled independently and is unaffected by the blocks/suckers. Note that when all suckers are controlled by a single agent, it is possible to learn the contraction patterns for coordinated crawling without assuming a CPG, as shown in Ref. [13].
Suckers are labeled by their ID $n$, where $n=1,\dots,N_{s}$ and $N_{s}$ is the number of suckers. The $n$-th spring is defined as the one to the right of the $n$-th sucker, and its equilibrium position is given by the CPG:
@@
where $k=2\pi/N$ with $N$ the periodicity, $\omega$ the angular velocity, $a$ a constant and uniform amplitude, and $x_{0}$ the natural length of the spring. This latter parameter fixes the spatial scaling of the model. Here we fix $N=N_{s}$, so that the wavelength $\lambda$ corresponds to the length of the crawler $L$. Arbitrary values of the periodicity are considered in Appendix C, where we show how $N$ alters the importance of each sucker when a crawling strategy is played. To mathematically define the possibility of adhesion by the suckers, we assume that each sucker has a distinct linear viscous coefficient, $\zeta_{n}$. As detailed in [15], without coordinated adhesion of each sucker/block, the crawler just oscillates along its center of mass and the average velocity is zero. We assume that each friction coefficient can have an identical constant value $\zeta_{0}$ (no adhesion) or $\zeta_{n}\rightarrow\infty$ (the $nth$ sucker adheres). The choice of infinite adhesion is one of the ingredients distinguishing our model from [15] where calculations are conducted under the assumption that individual suckers only slightly perturb their friction coefficient. In Appendix A we show the details of the physical model and its parametrization and scaling. In the following, to compare crawlers of different lengths, we will express velocities in normalized units $v\rightarrow v/x_{0}$ so that the results are independent of spatial scaling $x_{0}$.
@@
II.1.1 Expected qualitative behavior from the idealized continuous infinite crawler
In Ref.[15] the authors find an analytical solution for the velocity of the center of mass assuming that the 1D crawler is infinite and approximable to the continuous limit ($x_{0}\rightarrow 0$), and that the adhesion is a perturbation. For the sake of clarity we summarize here their results. The perturbation is taken in the form of a traveling wave pulse where only a single sucker can increase its viscous friction at a time, and with the same periodicity and frequency of the CPG (eq. 1),
$\zeta(s,t)=\zeta_{0}(1+\epsilon p(s,t))$ (2)
with $\epsilon$ a small parameter, and
$p(s,t)=\sum_{n=-\infty}^{\infty}2\pi\delta\left(\omega t-(ks-\beta-2n\pi)-\Theta\right)$ (3)
with $\Theta=\tan^{-1}\tilde{\omega}$ and $\tilde{\omega}=\zeta_{0}\omega/(k^{2}\kappa)$. The extra shift $\Theta$ to the CPG of eq. 1 comes from the solution to the unperturbed problem (see eq. 13, in Appendix B.1), that is, the suckers evolution without adhesion (no net motion of the crawler’s center of mass). With these assumptions, they find that the crawling velocity is proportional to
$\frac{v_{CM}}{x_{0}}\sim-\sqrt{\omega}\frac{\tilde{N}}{\sqrt{1+\tilde{N}^{4}}}\sin\beta$ (4)
with $\tilde{N}=\sqrt{\omega}N/(2\pi)$. Therefore, the optimal positioning of the pulse with respect to the traveling wave has a phase shift of $\beta-\Theta$, with $\beta=3/2\pi$. At this phase shift and for a given frequency $\omega$ of the CPG, a maximum is expected as a function of the spatial periodicity of the pulse, at
$N_{\max}=\frac{2\pi}{\sqrt{\omega}}$ (5)
Our model differs from this analytically tractable model in several ways: our crawler is discrete and finite; the head and tail play an important role as discussed below; suckers do not apply a small perturbation to the viscous resistance but rather adhere vigorously to the substrate; finally multiple suckers may adhere at any given moment. In Appendix B we discuss and compare extensively our model and simulator to Ref. [15]. Although we cannot expect quantitative correspondence of our crawler with the analytical model we do find an optimum as a function of the periodicity $N$, in qualitative agreement with Ref. [15].
II.2 Learning architectures
@@
Architecture $\epsilon$ $lr$ minimum training steps $\times$ episode and per agent
@@
Distributed (standard) $\lesssim 10^{-2}$ $\sim 10^{-3}$ minimum $\sim 50$k steps111difficult to converge and $N_{s}$ dependent
@@
Centralized (large33310 or 15 suckers per control center. CC) $\sim 10^{-2}$ $\gtrsim 10^{-1}$ $\gtrsim 100$k – $\sim 1000$k steps, about half for hive
Fully centralized (single CC) $\sim 10^{-1}$ robust up to $\gtrsim 10^{-1}$ $\gtrsim 100$k – $\sim 1000$k steps
@@
Table 1: Approximate value of the exploration parameters used for the various architectures, considering different sizes and number of CCs. The precise value is case specific and was tuned manually by considering the trade-off between satisfactory levels of exploration and convergence.
@@
In the Reinforcement Learning (RL) paradigm, the learning problem is framed in terms of Environment and Agent. The agent, i.e. the learner, interacts continuously with the environment selecting at each time step an action (stick or not stick to the substrate) to which the environment responds with a numerical reward and a representation of its state [16]. When treating the 1D crawler model as an RL problem, since the task is to achieve unidirectional and fast motion, the reward at each time step is given by the instantaneous center of mass velocity, $R_{t}=v_{CM}(t)$ when the velocity is positive, and $R_{t}=-1$ if $v_{CM}(t)<0$. After several trials, we found this choice to be effective to promote policies maximizing the crawling speed. In order to keep the problem as simple as possible and minimize the requirements on proprioception, we use binary states associated to the spring’s tension: 1 = elongated; 0 = compressed. As previously anticipated, actions are also binary and related to each sucker being adhering or not: $\zeta_{n}=\infty$ or $\zeta_{n}=\zeta_{0}$, respectively. Learning can be achieved via several algorithms. We here use standard tabular Q-learning, where a quality matrix, or Q-matrix, is learned with an iterative process. The Q-matrix is used to build a (deterministic) policy, consisting in selecting at each state the action that maximizes the empirical return which is an estimate of the expected cumulative reward when following the current policy. For a given agent its Q-matrix follows the update rule
$\begin{split}Q_{t+1}(S,A)=&Q(S_{t},A_{t})+\\ &\alpha_{t}\left[R_{t+1}+\gamma\max_{a}Q_{t}(S_{t+1},a)-Q(S_{t},A_{t})\right]\end{split}$ (6)
where $\alpha_{t}$ is the learning rate at integration step $t$ and $\gamma$ the discount. The learning parameters $\alpha$, $\gamma$ are discussed in the next section. Therefore, the dimensionality of the Q-matrix is given by $|\mathcal{S}|\times|\mathcal{A}|$, with $\mathcal{S}$ and $\mathcal{A}$ the ensembles of states and action, respectively. As illustrated in fig. 1, agency can be defined by the learning architecture in several ways. Each architecture carries a specific definition of states and actions which dictates the dimensionality of the Q-matrix, hence the computational cost. Finally, we assume that each agent can act on temporal scales comparable to the CPG oscillation frequency, so that responses are instantaneous. To investigate the role of centralization vs distribution, we explore two families of learning architectures, which we illustrate next.
II.2.1 Distributed control

### v3

@@ I Introduction
Cephalopods, such as octopus, squid and cuttlefish, have evolved a complex nervous system whose size is comparable to that of vertebrates and endows them with impressive cognitive capabilities. They do have a central brain; however most of their nervous system is distributed throughout their arms which are covered in hundreds of suckers that can sense the environment as well as adhere to substrates. Like all soft-bodied animals, these organisms have to coordinate many degrees of freedom to achieve locomotion. For example, octopuses translocate on surfaces by leveraging their taste-by-touch sense, whereby suckers adhere to surfaces and sense them chemically upon contact [1]. The sensory information acquired by the periphery guides suckers’ adhesion, yielding displacement of the organism along the surface. However, to what extent locomotion is controlled by the central nervous system vs by the ganglia distributed along the arms is not known [34, 31]. Intriguingly, coordination in octopus may or may not depend on precise proprioceptive feedback, since the molecular mechanisms for proprioception and to what extent the brain stores positional information is not understood [17, 61].
This question is part of a large body of literature in neuromechanics that broadly asks how do animals and robots control various kinds of distributed locomotion to achieve specific goals including energy efficiency, maneuverability, stability, speed (see e.g. [11, 39] and references therein). Approaches that range from a more centralized [23] to a purely distributed control [54] have been put forward for various locomotion mechanisms in bio-inspired robotics, reviewed e.g. in [46]. Whether centralization increases locomotion speed is an open question (reviewed e.g. in [38]). On the one hand, centralization likely requires additional processing which may be too slow to be implemented at fast speed. On the other hand, the need for added stability at high speed may require centralized control. The combination of CPG and local feedback was found beneficial for adaptation and versatility for amphibious centipedes [59], undulatory patterns in swimming and multi-leg terrestrial locomotion [58] and quadruped locomotion [50].
@@
Here we focus on adhesion-driven crawling: we draw inspiration from octopus and focus on a toy model to ask what are the pros and cons of centralization vs distribution. The crawler is modeled as a collection of suckers distributed along a straight line and connected by springs, and endowed with a rudimentary form of proprioception (extension vs elongation of the springs) and a rudimentary form of control of the suckers (adhere vs not adhere). Contraction of the springs is not controlled by the crawler but rather follows a stereotyped wave established by a central pattern generator (CPG). If suckers were to adhere at random times, the crawler center of mass would be unable to move on average: is it possible to learn a pattern of adhesion that enables net translocation, purely from trial and error?
Learning is an underexplored aspect of crawling that is relevant to develop robotic crawlers in unpredictable environments, for example on irregular terrain or under other environmental perturbations. In these cases, programming the robot to crawl is unfeasible and learning from trial and error becomes a viable option.
To answer this question we develop a simulator of the crawler and let it learn to translocate by trial and error using the Q-learning Reinforcement Learning algorithm. We take the approach of fixing a widely used model of coupled oscillators, or template [14], where each unit has the ability to adhere. The mathematical model is similar to a large body of literature on the neuromechanics of crawling largely developed for larvae, worms and insects (e.g. [4, 29, 40] and references within). Our crawler does not control muscle contraction, but only adhesion with the substrate. A main difference with most literature (but see [36]) is that we do not program the crawler, but let it learn through the use of reinforcement learning, whereby we set the goal of translocating and the agent learns to crawl purely from trial and error. This approach allows us to assess how easy it is for a learning architecture to develop an efficient crawling, and how computationally demanding learning is. Efficiency is defined as the combination of fast translocation and robustness to failure of individual suckers. Alternatively, one may target a specific dynamics and learn how to formulate a mathematical model that reproduces a specific attractor in the phase space [24]. We ask whether the sole control of adhesion is sufficient to learn crawling with a distributed vs centralized architecture. Our results complement previous literature considering patterns of contraction –rather than adhesion– in fully centralized learning of crawling [36] and fully distributed learning of swimming [18], as well as works where crawling emerges from prescribed control patterns, rather than learning (see e.g. [51, 44]), and where the redundancies of preprogrammed gait design lead alone to locomotion robustness, for instance over rough terrains [9].
@@
We compare the learnability, performance and robustness of centralized vs distributed learning architectures. Despite the rudimentary proprioception and control, we find that the sole existence of a common reward –here the speed of the center of mass– is sufficient for independent suckers to learn crawling. When all suckers are simultaneously controlled by a single agent, crawling is both faster and more robust to failure, partially overcoming the crawler’s poor sensorimotor ability. However, the computational cost of the central control scales exponentially with the number of suckers and springs. We then introduce a shallow hierarchical architecture where control is centralized in few control centers. We find that this partial centralization achieves nearly optimal performance and robustness with a limited increase in computational burden.
In our setting, crawling is achieved by “surfing” the wave generated by the central pattern generator. The existence of a central pattern generator is commonly assumed and has been demonstrated in various systems (see e.g. [16]). Note however that in the absence of a central pattern generator, crawling may still be achieved by proprioceptive feedback alone [57, 7, 41], or by leveraging synchronization of motor control with some natural frequency of the system (for example in the context of human inspired biped locomotion via imitation learning [37], or limb oscillation via experimental analysis [20], or for flapping wings in flight [32]).
@@
We hope that our work will serve as a benchmark to explore the optimal level of centralization in more realistic scenarios for natural and artificial crawling, where constraints on speed, robustness and computational costs will all concur to define optimality.
@@ II Model and learning
[…]
@@ II.1 The 1D crawler model
The model, illustrated in fig. 1, is strongly inspired by the one proposed in [51] and considers the crawler as a series of suckers (blocks) connected by springs along the body axis. The crawler is placed on a flat substrate and its dynamics is dictated by the springs and the interaction with the substrate:
$split m\ddot{x}_{n}(t)=-\zeta_{n}\dot{x}_{n}(t)+\kappa\left(x_{n+1}(t)-x_{n}(t)-l_{n}(t)\right)&\\ -\kappa\left(x_{n}(t)-x_{n-1}(t)-l_{n-1}(t)\right)$ (1)
where $x_{n}$ is the position of the $n$th sucker, with $n=1,\dots,N_{s}$; $N_{s}$ is the number of suckers; $m$ is the mass of each sucker, $\zeta_{n}$ its friction coefficient and $\kappa$ the spring constant; $l_{n}$ is the rest length of the $n$-th spring and it oscillates in time as prescribed by the CPG as described below. We assume friction is proportional to the sliding velocity of the suckers, modeling viscous friction with a wet substrate. Each sucker can increase its viscous coefficient $\zeta_{n}$ by adhering with the substrate. For simplicity, we assume that each friction coefficient can select only two values: a constant value $\zeta_{0}$ (no adhesion) or $\zeta_{n}\rightarrow\infty$ (the nthnth sucker adheres). The choice of infinite adhesion is one of the ingredients distinguishing our model from [51] where calculations are conducted under the assumption that individual suckers only slightly perturb their friction coefficient.
For the tail, $n=0$ , and the head, $n=N_{s}$, the dynamics is driven by a single adjacent spring: {align} m¨x_0(t) = -ζ_0˙x_0(t) + κ(x_1(t)-x_0(t)-l_0(t))
m¨x_N_s(t) =-ζ_0˙x_N_s(t) -κ(x_N_s(t)-x_N_s-1(t)-l_N_s-1(t))
II.2 Central pattern generator (CPG)
We assume an elongation-contraction wave is generated by a CPG and propagates from the tail to the head of the crawler:
@@ II.2 Central pattern generator (CPG)
where $l_{n}$ is the rest length of the $n$-th spring, defined as the one to the right of the $n$-th sucker; $k=2\pi/N$ with $N$ the periodicity, $\omega$ the angular velocity, $a$ a constant and uniform amplitude, and $x_{0}$ the natural length of the unperturbed spring. The unperturbed rest length of the springs, $x0$, fixes the spatial scaling of the model. The spatial periodicity $N$ in the CPG oscillation is in principle an arbitrary integer. For simplicity, we fix $N=N_{s}$, so that the wavelength $\lambda=x_{0}N$ corresponds to the length of the crawler $L$. Arbitrary values of the periodicity are considered in Appendix B, where we show how $N$ alters the importance of each sucker when a crawling strategy is played. Note that when all suckers are controlled by a single agent, it is possible to learn the contraction patterns for coordinated crawling without assuming a CPG, as shown in Ref. [36]. We assume the CPG is unaffected by the suckers, relevant to real locomotion in the case where proprioceptive feedback has low gain [53].
II.3 Scaling and parametrization
[…]
$\cases{\dot{\hfil}}{\bar{x}}_{n}(t)=B\left(\bar{x}_{n+1}(t)+\bar{x}_{n-1}(t)-2\bar{x}_{n}(t)-\bar{l}_{n}(t)+\bar{l}_{n-1}(t)\right)\\ \dot{\bar{x}}_{0}(t)=B\left(\bar{x}_{1}(t)-\bar{x}_{0}(t)-\bar{l}_{0}(t)\right)\\ \dot{\bar{x}}_{N_{s}}(t)=B\left(\bar{x}_{N_{s}-1}(t)-\bar{x}_{N_{s}}(t)+\bar{l}_{N_{s}-1}(t)\right)\\ \bar{l}_{n}=1+\bar{a}\sin(\bar{t}-2\pi n/N)$ (3)
where the nondimensional group $B_{n}=\kappa/(\zeta_{n}\omega)$ is finite for suckers that do not adhere ($B_{n}=B_{0}$) and it vanishes for suckers that adhere ($B_{n}=0$ i.e. when the $n$-th sucker adheres its velocity drops instantaneously to zero, $\dot{\bar{x}}_{n}=0$). The problem is thus dictated by 4 nondimensional parameters:
{B0=κ/(ζ0ω)=10N=Ns except for fig. 10
Ns
=L/x0 varying
a
/x0
=0.25
[…]
@@ II.3 Scaling and parametrization
II.3.1 Expected qualitative behavior from the idealized continuous infinite crawler
In Ref. [51] the authors find an analytical solution for the velocity of the center of mass of a continuous ($N_{s}\rightarrow\infty$, $x_{0}\rightarrow 0$ and $Nx_{0}=\lambda$) one dimensional infinitely long crawler, assuming individual blocks can only perturb slightly the friction coefficient (rather than fully adhere like in our case). For the sake of clarity we summarize here their results. The $n$-th friction coefficient is perturbed according to:
$\zeta_{n}(t)=\zeta_{0}(1+\epsilon p(s,t))$ (5)
[…]
$v\propto\frac{\tilde{\omega}}{\sqrt{1+\tilde{\omega}^{4}}}$ (6)
where $\tilde{\omega}=\omega\zeta_{0}N^{2}/(4\pi^{2}\kappa)$. Note that this equation has a maximum at $\tilde{\omega}=1$, which occurs when
$N_{\text{max}}=2\pi\sqrt{\kappa/(\omega\zeta_{0})}.$
This maximum is achieved when the wavelength of the CPG, $\lambda=Nx_{0}$, matches a diffusion lengthscale $\lambda_{D}=\sqrt{D/\omega}$ where $D=kx_{0}^{2}/\zeta_{0}$ can be derived analytically as the continuous model reduces to an effective diffusive equation. While our finite crawler cannot be treated analytically as in Ref. [51], the diffusive lengthscale is still dimensionally correct, and we expect that qualitatively the crawler finds a maximum speed when $N_{\text{max}}\sim\sqrt{\kappa/(\omega\zeta_{0})}$. We do indeed observe that the crawler translocates with a speed that depends on the number of suckers, with a maximum at an intermediate location with about $N=12$ suckers (Fig. 3). The location of the maximum varies with the parameters as prescribed by the scaling above (data not shown), suggesting that the physics of this optimal condition is robust to details of the model. In Appendix A we discuss and compare extensively our model and simulator to Ref. [51].
Note that living crawlers may evolve to match this constraint by e.g. developing specific secretions that will modify their friction coefficient $\zeta_{0}$, or the frequency $\omega$ or wavelength $\lambda$ of their CPG, or the stiffness of their muscles $\kappa$. In the following, we do not focus on the mechanical parameters that define the crawler, but rather on how to learn the controls that enable it to move.
II.4 Learning architectures
@@
Architecture $\epsilon$ $\alpha$ minimum training steps $\times$ episode
@@
Distributed (standard) $\lesssim 10^{-2}$ $\sim 10^{-3}$ $\sim 50$k steps111difficult to converge and dependent on the number of suckers (agents)
@@
Centralized (large33310 or 15 suckers per control center. CC) $\sim 10^{-2}$ $\gtrsim 10^{-1}$ from $\gtrsim 100$k to $\sim 1000$k steps, about half for hive
Fully centralized (single CC) $\sim 10^{-1}$ robust up to $\gtrsim 10^{-1}$ from $\gtrsim 100$k to $\sim 1000$k steps
@@
Table 1: Approximate value of the exploration parameters used for the various architectures trained, including the number of integration steps per episode needed to obtain convergence (consider also fig. 2). The precise value is case specific and was tuned manually by considering the trade-off between satisfactory levels of exploration and convergence.
@@
In the Reinforcement Learning (RL) paradigm, the learning problem is framed in terms of Environment and Agent. The agent, i.e. the learner, interacts continuously with the environment selecting at each time step an action (stick or not stick to the substrate) to which the environment responds with a numerical reward and a representation of its state [49]. When treating the 1D crawler model as an RL problem, since the task is to achieve unidirectional and fast motion, the reward at each time step is given by the instantaneous center of mass velocity, $R_{t}=v_{CM}(t)$ when the velocity is positive, and $R_{t}=-1$ if $v_{CM}(t)<0$. After several trials, we found this choice to be effective to promote policies maximizing the crawling speed. In order to keep the problem as simple as possible and minimize the requirements on proprioception, we use binary states associated to the spring’s tension: 1 = elongated; 0 = compressed. As previously mentioned, actions are also binary and related to each sucker being adhering or not: $\zeta_{n}=\infty$ or $\zeta_{n}=\zeta_{0}$, respectively. Note that the instantaneous reward $R_{t}$ is a random variable, as the net speed of the center of mass is the result of the coupled spring dynamics which is only partially sensed and controlled by the crawler. We thus seek to maximize the cumulative translocation over many steps in the future, as usual in RL. To do this, we introduce the long term reward $G_{t}$, defined as $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}$ where $R_{t+k+1}$ is the instantaneous reward $k+1$ steps in the future with respect to $t$. We use the standard formulation with discount rate $\gamma<1$, which is a parameter discounting rewards further in the future ($\gamma=0$ truncates the sum to its first term i.e. $G_{t}=R_{t+1}$ is the instantaneous reward; whereas $\gamma=1$ weighs each contribution equally for all times). We use $\gamma=0.99$, corresponding to truncating the sum to an effective horizon of $100$ steps. For continuing tasks, another possible formulation of the reward is that of average reward [49, 33], which does not require discount and where the reward becomes also an estimate. Preliminary results however do not indicate any visible benefits of this formulation in our problem.
Learning seeks to find a policy of actions, $\pi(a|s)$, i.e. a rule that prescribes how to choose actions $a$ at each given state $s$ that maximizes the expected future reward. Optimal policies can be searched via several algorithms. We here use standard tabular Q-learning, where a quality matrix, or Q-matrix, is learned with an iterative process. The Q-matrix is the expected cumulative reward conditioned to being in state $s$ and choosing action $a$, i.e. $Q(s,a)=E(G_{t}|S_{t}=s,A_{t}=a)$, while continuing with policy $\pi$. $Q$ is estimated iteratively and each estimate is used to build a policy consisting in selecting at each state the action that maximizes the empirical return with highest probability $(1-\epsilon_{t})$ and choosing a random action with a small probability $\epsilon_{t}/|\cal{A}|$. Thus for a given agent its Q-matrix follows the update rule
[…]
II.4.1 Distributed control